OpenStudy (logan13):

If the diameter of the circle above is 12 mm, what is the area of the circle? 6 mm2 36 mm2 144 mm2 12 mm2

OpenStudy (unanimoose):

Ok so before you keep posting the questions to your test let me walk this out for you so that you can do it by yourself without having others answer it for you.

OpenStudy (logan13):

@AlexandervonHumboldt2

OpenStudy (unanimoose):

\[A = r^2\pi\]

OpenStudy (alexandervonhumboldt2):

radius=1/2 diameter area=radius^2*3.14

OpenStudy (logan13):

not a test

OpenStudy (unanimoose):

The radius is needed to do this. The radius is HALF the diameter.

OpenStudy (unanimoose):

If it's not a test then what is it?

OpenStudy (logan13):

studyisland

OpenStudy (unanimoose):

But anyway, the diamater is 12 here so to find the radius we need to do: 12/2

OpenStudy (unanimoose):

If you're just going to look for answers on here then you may as well stop using studyisland because it's not doing anything for you.

OpenStudy (unanimoose):

So anyway: 12/2 = 6 Now we can plug that in: \[A = 6^2\pi\]

OpenStudy (unanimoose):

So now we can square the radius and you will have you solution, and for most options you won't need to apply pi. \[A = 36\pi\]